OpenStudy (anonymous):

new question :) (-9x ÷ (x + 3) - (-3 ÷ (x +3)

OpenStudy (pfenn1):

\[\frac{-9x}{x+3}-\frac{-3}{x+3}\]

OpenStudy (pfenn1):

\[\frac{-9x+3}{x+3}\]

OpenStudy (pfenn1):

Can you simplify it further?